OpenStudy (anonymous):

4^2 over 8x^2-8x = 2 over x

OpenStudy (bahrom7893):

\[\frac{4^2}{8x^2-8x}=\frac{2}{x}\]?

OpenStudy (anonymous):

yes but the 4^2 is 4x^2 , sorry typo and i cross mult and got 2(8x^2-8x)=4x^2 times x and then i got 16x^2-16x=4x^3

OpenStudy (bahrom7893):

okay so: \[\frac{4x^2}{8x^2-8x}=\frac{2}{x}\]

OpenStudy (anonymous):

am i right so far? and yes thats it (:

OpenStudy (anonymous):

am i right so far? and yes thats it (:

OpenStudy (bahrom7893):

\[4x^3=16x^2-16x\]\[4x^3-16x^2+16x=0\]\[4x(x^2-4x+4)=0\]

OpenStudy (bahrom7893):

Can you continue from here? All I did was move everything to one side and pull out a 4x.

OpenStudy (anonymous):

why isnt the 4x^3 not negative?

OpenStudy (bahrom7893):

because it was already on that side. For example if we wanted to move 3 over in this case: x+3=-7 We would do: (x + 3) - 3 = (-7) - 3 x = -7 - 3

OpenStudy (anonymous):

oh okay i see so now i can solve that equation using the quad. formula right?

OpenStudy (bahrom7893):

yup.

OpenStudy (anonymous):

okay thank you (:

OpenStudy (bahrom7893):

And you know that x is either 0, or whatever the solutions to: \[x^2-4x+4=0\]are

OpenStudy (anonymous):

when i put it in the quad form i dont use their exponents right?

OpenStudy (bahrom7893):

well actually that equation is fairly straightforward to solve: \[x^2-4x+4=x^2-2*2*x+2^2=(x-2)^2\]

OpenStudy (bahrom7893):

so your answer would be x=0, x=2 (with x=2 being a repeated root)

OpenStudy (anonymous):

thats not one of my options ): @bahrom7893

OpenStudy (anonymous):

thats not one of my options ): @bahrom7893

OpenStudy (bahrom7893):

0 and 2 are not there?

OpenStudy (anonymous):

no there is x=2 and x=-2 , x=2, x=16 and x=-16, x=-16

OpenStudy (bahrom7893):

ohh x=2 x cannot be 0 because of 2/x

OpenStudy (anonymous):

okay thank you (: